OpenStudy (anonymous):

The measure of angle A to the nearest degree is 18 19 20

OpenStudy (anonymous):



OpenStudy (anonymous):

sinA-3/9 =1/3

OpenStudy (anonymous):

x = arcsin(1/3) = 19.47º

OpenStudy (stormfire1):

It should be 19.47...so 19.

OpenStudy (anonymous):

thx